OpenStudy (anonymous):

From the top of a cliff, the angle of depression to a boat out in the water is 28 degrees. If the boat is 750 feet away from the bottom of the cliff, how high is the cliff off the ground?

OpenStudy (anonymous):

Angle of depression equals the angle of the elevation from the boat's point of view.\[\text{height}=750 \tan (28 {}^{\circ})=398.8 \text{ feet} \]

OpenStudy (ash2326):

 we know perpendicular /base=sin theta theta=28 p=750 b=? so sin 28=750/b b=750/sin(28) b=1597.7844 meters that's the height of the cliff